How many discs of one-inch diameter can you fit in a rectangle of size 5 inches by 3.6 inches?  (In a sense, the capacity of the rectangle with respect to the discs)

(This is a two-dimensional problem - you cannot stack discs)
Answer #1

Total = 5+4+5+4 = 18

Arrangement will be like:

Area of disc = pi*((d/2)^2 = 3.14*(0.5)^2 = 0.785

Area of rectangle = l*b = 5*3.6 = 18

Maximum Number of disc possible = Area of rectangle/Area of disc = 18/0.785 = 22.9

So, 18 is possible basis hit and trial